OpenStudy (anonymous):

limit as x approaches 0. (e^(2x) -1) / ( 1- e^X)

ganeshie8 (ganeshie8):

see if u can factor numerator

ganeshie8 (ganeshie8):

e^(2x) -1 = (e^x)^2 - 1^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what should i do next ??

ganeshie8 (ganeshie8):

do u knw this identity ? \(a^2 - b^2 = (a+b)(a-b)\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x\to 0} ~\frac{e^{2x}-1}{1-e^x}\) \(\large \lim \limits_{x\to 0} ~\frac{(e^{x})^2-1^2}{1-e^x}\)

ganeshie8 (ganeshie8):

apply that identity to numerator

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x\to 0} ~\frac{e^{2x}-1}{1-e^x}\) \(\large \lim \limits_{x\to 0} ~\frac{(e^{x})^2-1^2}{1-e^x}\) \(\large \lim \limits_{x\to 0} ~\frac{(e^x+1)(e^x-1)}{1-e^x}\)

ganeshie8 (ganeshie8):

can u do the rest ha ?

OpenStudy (anonymous):

so now we can eliminate 1-e^X right ?

ganeshie8 (ganeshie8):

yes thats the reason we factored :)

ganeshie8 (ganeshie8):

do it, and let me knw the final answer u get

OpenStudy (anonymous):

is it 2

ganeshie8 (ganeshie8):

close, but no. try again

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x\to 0} ~\frac{e^{2x}-1}{1-e^x}\) \(\large \lim \limits_{x\to 0} ~\frac{(e^{x})^2-1^2}{1-e^x}\) \(\large \lim \limits_{x\to 0} ~\frac{(e^x+1)(e^x-1)}{1-e^x}\) \(\large \color{red}{-}\lim \limits_{x\to 0} ~\frac{(e^x+1)(e^x-1)}{e^x-1}\) \(\large \color{red}{-}\lim \limits_{x\to 0} ~e^x+1\) take the limit now \(\large \color{red}{-} (e^0 + 1)\) \(\large \color{red}{-} (1 + 1)\) \(\large \color{red}{-} (2)\) \(\large -2\)

OpenStudy (anonymous):

thanks a lot :)

ganeshie8 (ganeshie8):

np.. u wlc :)